Suppose you graphS i n open parentheses theta subscript 1 close parentheses Vs. S i n open parentheses theta subscript 2 close parentheses and get a slope m = 0.413. If the second material is air (n2 = 1.0003), use equation 2 in the manual to calculate n1.

n1 = 1

n1 = 2.42

n1 = 1.46

n1 = 1.47

n1 = 1.49

n1 = 0.413
Answer #1

Here, According to snells law

=> n1 * sin(theta1) = n2 * sin(theta2)

=> n1 * sin(theta1) = 1.0003 * sin(theta2)

=>   n1 = 1.0003 * sin(theta2)/sin(theta1)

             =   1.0003 * 1/0.413

              =   2.42

=>   n1 = 2.42